# Beyond Cosine Similarity: Magnitude-Aware CLIP for No-Reference Image Quality Assessment

### v2

@@ Introduction
Recently, significant progress has been made in learning-based NR-IQA models. However, most existing approaches rely heavily on supervised training with specific IQA datasets, often overfitting to dataset-specific distortions or content. This generalization bottleneck severely limits their applicability in real-world systems. The emergence of large-scale vision-language models such as CLIP (radford2021learning) offers a promising alternative. Trained on hundreds of millions of image–text pairs, CLIP demonstrates remarkable generalization and semantic understanding in a zero-shot manner. Recent works have adapted CLIP for NR-IQA by leveraging its ability to compute the cosine similarity between image embedding and quality-descriptive textual prompts (e.g., “a good photo” vs.“a bad photo”) (wang2023exploring). This prompt-based CLIP-IQA technique has shown surprising effectiveness on standard IQA benchmarks without any fine-tuning, providing a compelling direction for generic, training-free quality assessment.
@@
Figure 2: Overview of the Proposed Magnitude-Aware CLIP IQA Framework. Given an input image, we extract its CLIP image embedding and compute two quality signals: (1) $Q_{\text{sim}}$, the image semantic similarity with text prompts, and (2) $Q_{\text{mag}}$, a magnitude-based score obtained via Box-Cox transformation for statistical normalization. To adaptively balance these complementary cues, we compute a confidence discrepancy and generate softmax-based fusion weights, producing the final quality prediction $Q$.
@@
Early no-reference IQA approaches relied on handcrafted features that captured natural scene statistics (NSS), with descriptors derived from spatial (mittal2012making), wavelet (moorthy2010two), and DCT (saad2012blind) domains. Psychovisual models inspired by the free-energy principle of the human visual system further enriched the modeling of perceptual degradation (gu2014using; zhai2011psychovisual). With the advent of deep learning, CNN-based models became prevalent. Early works such as Kang et al. (kang2014convolutional) learned quality-aware representations directly from image patches, and subsequent extensions introduced multi-task training (kang2015simultaneous) to jointly predict quality and distortion types. Later models such as DIQaM (bosse2017deep), FPR (chen2020no), and GraphIQA (sun2022graphiqa) further advanced performance by modeling spatial dependencies and incorporating relational reasoning. Transformer-based models have also been explored due to their strong modeling capacity and flexibility. You et al. (you2021transformer) introduced a Transformer backbone for NR-IQA that benefits from features pretrained on large-scale classification tasks. Other approaches leverage perceptual priors from image restoration networks (lin2018hallucinated; chen2022no). However, these models struggle to generalize due to limited data and domain-specific bias. To address this, recent work incorporates auxiliary tasks (you2021transformer; lin2018hallucinated) and learning strategies such as meta-learning (zhu2020metaiqa), curriculum learning (wang2023hierarchical), and domain generalization (chen2021learning; chen2025monotonic). Despite progress, generalization in NR-IQA remains an open problem.
@@
Vision-language models provide a new paradigm for NR-IQA. Wang et al. (wang2023exploring) first leveraged CLIP (radford2021learning) to compute the similarity between distorted images and antonym-paired textual prompts. Follow-up work such as IPCE (peng2024aigc) mapped cosine similarity scores to discrete quality levels using hand-crafted prompts, while CLIP-AGIQA (tang2024clip) learned prompt tokens for six quality classes and concatenated them with image features for quality regression. In parallel, benchmarks such as Q-Bench (wu2023qbench) and DepictQA (you2024descriptive) evaluated the IQA capability of large multimodal models (LMMs) on low-level visual perception. Zhu et al. (zhu20242afc) examined 2AFC-style prompting for preference judgment, and Q-Align (wu2023align) proposed rating-level prompts to elicit more consistent predictions. Further improvements have been achieved by leveraging pre-trained LMMs and high-quality instruction datasets (wu2024towards; zhu2024adaptive). However, fine-tuning these models on IQA tasks often results in catastrophic forgetting (luo2023empirical), weakening performance on other domains. This motivates the need to explore training-free strategies that better preserve the generality of pre-trained vision-language models.
@@
To mitigate the semantic bias inherent in raw CLIP embeddings, we introduce a statistical normalization approach based on the Box-Cox transformation (box1964analysis), which is a classical power transform that stabilizes variance and reduces distributional skewness. Unlike cosine similarity, which discards magnitude information via $\ell_{2}$ normalization, our goal is to retain and standardize this information for quality prediction. Given the image-level feature embedding ${F}_{{img}}\in\mathbb{R}^{D}$, we first take its element-wise absolute value to remove the polarity and retain only activation strength:
@@
We evaluate our method on a wide range of datasets to verify its generalization and robustness. These include: (1) Synthetic distortion datasets: CSIQ (larson2010most), TID2013 (ponomarenko2015image), and KADID-10k (lin2019kadid), containing various artificially generated distortions. (2) Authentic distortion datasets: CLIVE (ghadiyaram2015massive), KonIQ (hosu2020koniq), and SPAQ (fang2020perceptual), reflecting real-world degradations from mobile photography. (3) Image restoration (IR) datasets: PIPAL (jinjin2020pipal), featuring restored images from SR/denoising/deblurring pipelines. (4) AIGC quality datasets: AGIQA-1k (zhang2023perceptual) and AGIQA-3k (li2023agiqa), designed for quality assessment of AI-generated content. All evaluations follow the standard protocol, where Spearman’s Rank Correlation Coefficient (SRCC) and Pearson Linear Correlation Coefficient (PLCC) are reported.
@@
We compare our method with two categories of existing IQA approaches: (1) Opinion-Unaware (OU) Methods. These methods do not require subjective opinion scores for training and are typically used in a zero-shot manner. We include NIQE (mittal2012making), QAC (xue2013learning), PIQE (venkatanath2015blind), LPSI (wu2015highly), dipIQ (ma2017dipiq), SNP-NIQE (liu2019unsupervised), NPQI (liu2020blind), CLIPIQA (wang2023exploring), ContentSep (babu2023no), and MDFS (ni2024opinion) in this category. Since our method is also a zero-shot approach that requires no training on opinion scores, this group serves as the most appropriate baseline for a fair comparison. (2) Learning-based Methods. These methods rely on training with human-annotated quality scores and are typically optimized for specific IQA datasets. We include Re-IQA (saha2023re), ARNIQA (agnolucci2024arniqa), CLIP-IQA+ (wang2023exploring), and GRepQ (srinath2024learning). For a comprehensive evaluation of generalization capability, we compare with these learning-based methods trained on the training split of each testing dataset.

### v3

@@ Introduction
Recently, significant progress has been made in learning-based NR-IQA models. However, most existing approaches rely heavily on supervised training with specific IQA datasets, often overfitting to dataset-specific distortions or content. This generalization bottleneck severely limits their applicability in real-world systems. The emergence of large-scale vision-language models such as CLIP (Radford et al. 2021) offers a promising alternative. Trained on hundreds of millions of image–text pairs, CLIP demonstrates remarkable generalization and semantic understanding in a zero-shot manner. Recent works have adapted CLIP for NR-IQA by leveraging its ability to compute the cosine similarity between image embedding and quality-descriptive textual prompts (e.g., “a good photo” vs.“a bad photo”) (Wang et al. 2023a). This prompt-based CLIP-IQA technique has shown surprising effectiveness on standard IQA benchmarks without any fine-tuning, providing a compelling direction for generic, training-free quality assessment.
@@
Figure 2: Overview of the Proposed Magnitude-Aware CLIP IQA Framework. Given an input image, we extract its CLIP image embedding and compute two quality signals: (1) $Q_{\text{sim}}$, the image semantic similarity with text prompts, and (2) $Q_{\text{mag}}$, a magnitude-based score obtained via Box-Cox transformation for statistical normalization. To balance these complementary cues, we adopt a confidence discrepancy and generate softmax-based fusion weights, producing the final quality prediction $Q$.
@@
Early no-reference IQA approaches relied on handcrafted features that captured natural scene statistics (NSS), with descriptors derived from spatial (Mittal et al. 2012), wavelet (Moorthy and Bovik 2010), and DCT (Saad et al. 2012) domains. Psychovisual models inspired by the free-energy principle of the human visual system further enriched the modeling of perceptual degradation (Gu et al. 2014; Zhai et al. 2011). With the advent of deep learning, CNN-based models became prevalent. Early works such as Kang et al. (Kang et al. 2014) learned quality-aware representations directly from image patches, and subsequent extensions introduced multi-task training (Kang et al. 2015) to jointly predict quality and distortion types. Later models such as DIQaM (Bosse et al. 2017), FPR (Chen et al. 2020), and GraphIQA (Sun et al. 2022) further advanced performance by modeling spatial dependencies and incorporating relational reasoning. Transformer-based models have also been explored due to their strong modeling capacity and flexibility. You et al. (You and Korhonen 2021) introduced a Transformer backbone for NR-IQA that benefits from features pretrained on large-scale classification tasks. Other approaches leverage perceptual priors from image restoration networks (Lin and Wang 2018; Chen et al. 2022). However, these models struggle to generalize due to limited data and domain-specific bias. To address this, recent work incorporates auxiliary tasks (You and Korhonen 2021; Lin and Wang 2018) and learning strategies such as meta-learning (Zhu et al. 2020), curriculum learning (Wang et al. 2023b), and domain generalization (Chen et al. 2021, 2025). Despite progress, generalization in NR-IQA remains an open problem.
@@
Vision-language models provide a new paradigm for NR-IQA. Wang et al. (Wang et al. 2023a) first leveraged CLIP (Radford et al. 2021) to compute the similarity between distorted images and antonym-paired textual prompts. Follow-up work such as IPCE (Peng et al. 2024) mapped cosine similarity scores to discrete quality levels using hand-crafted prompts, while CLIP-AGIQA (Tang et al. 2024) learned prompt tokens for six quality classes and concatenated them with image features for quality regression. In parallel, benchmarks such as Q-Bench (Wu et al. 2023a) and DepictQA (You et al. 2024) evaluated the IQA capability of large multimodal models (LMMs) on low-level visual perception. Zhu et al. (Zhu et al. 2024a) examined 2AFC-style prompting for preference judgment, and Q-Align (Wu et al. 2023b) proposed rating-level prompts to elicit more consistent predictions. Further improvements have been achieved by leveraging pre-trained LMMs and high-quality instruction datasets (Wu et al. 2024; Zhu et al. 2024b). However, fine-tuning these models on IQA tasks often results in catastrophic forgetting (Luo et al. 2023), weakening performance on other domains. This motivates the need to explore training-free strategies that better preserve the generality of pre-trained vision-language models.
@@
To mitigate the semantic bias inherent in raw CLIP embeddings, we introduce a statistical normalization approach based on the Box-Cox transformation (Box and Cox 1964), which is a classical power transform that stabilizes variance and reduces distributional skewness. Unlike cosine similarity, which discards magnitude information via $\ell_{2}$ normalization, our goal is to retain and standardize this information for quality prediction. Given the image-level feature embedding ${F}_{{img}}\in\mathbb{R}^{D}$, we first take its element-wise absolute value to remove the polarity and retain only activation strength:
@@
We evaluate our method on a wide range of datasets to verify its generalization and robustness. These include: (1) Synthetic distortion datasets: CSIQ (Larson and Chandler 2010), TID2013 (Ponomarenko et al. 2015), and KADID-10k (Lin et al. 2019), containing various artificially generated distortions. (2) Authentic distortion datasets: CLIVE (Ghadiyaram and Bovik 2015), KonIQ (Hosu et al. 2020), and SPAQ (Fang et al. 2020), reflecting real-world degradations from mobile photography. (3) Image restoration (IR) datasets: PIPAL (Jinjin et al. 2020), featuring restored images from SR/denoising/deblurring pipelines. (4) AIGC quality datasets: AGIQA-1k (Zhang et al. 2023) and AGIQA-3k (Li et al. 2023), designed for quality assessment of AI-generated content. All evaluations follow the standard protocol, where Spearman’s Rank Correlation Coefficient (SRCC) and Pearson Linear Correlation Coefficient (PLCC) are reported.
@@
We compare our method with two categories of existing IQA approaches: (1) Opinion-Unaware (OU) Methods. These methods do not require subjective opinion scores for training and are typically used in a zero-shot manner. We include NIQE (Mittal et al. 2012), QAC (Xue et al. 2013), PIQE (Venkatanath et al. 2015), LPSI (Wu et al. 2015), dipIQ (Ma et al. 2017), SNP-NIQE (Liu et al. 2019), NPQI (Liu et al. 2020), CLIPIQA (Wang et al. 2023a), ContentSep (Babu et al. 2023), and MDFS (Ni et al. 2024) in this category. Since our method is also a zero-shot approach that requires no training on opinion scores, this group serves as the most appropriate baseline for a fair comparison. (2) Learning-based Methods. These methods rely on training with human-annotated quality scores and are typically optimized for specific IQA datasets. We include Re-IQA (Saha et al. 2023), ARNIQA (Agnolucci et al. 2024), CLIP-IQA+ (Wang et al. 2023a), and GRepQ (Srinath et al. 2024). For a comprehensive evaluation of generalization capability, we compare with these learning-based methods trained on the training split of each testing dataset.